evening shift 4 Jan 00


Subject: evening shift 4 Jan 00
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:38:59 CST


                Evening Shift 4 Jan 00
 
On Shift: Rick T, Marj
Also present: Rick K, Wah, Elliott, Jason H
Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
 
MI quad failure, est. as of 16:00 is 12 hrs
========================================================================
Runs ended this shift:

Run 15395 Cosmics
Run 15397 Cosmics

Tapes brought to FCC this shift: none
 

========================================================================
Accesses during shift:

Elliott and Jason, to replace power strip on ET system.
Marj and Rick T, to look at laser repeater signals

========================================================================
Overview of shift:

No beam, little activity. Ran cosmics most of the shift, trigger hung
often. Cycled laser crate repeater many times, KFREND many times.
Still a puzzle

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15395 16:00 MI quad failure...12 hr estimate. Continue
                       cosmic run to keep system alive
                17:05 Elliott and Jason make access to replace ET power strip
          
                17:20 Run manager checks in...he suggests we phone the
                       Mojove desert.
                17:40 Jason and Elliott out of access.

15395 19:15 End run so Rick K can add C13 trigger in preparation
                       for future tests

                19:30 Note that we have only one more bottle of Ar-ethane.
                       Have new bottles been ordered? Email sent to Dave N

15397 21:00 Start new cosmic run after Rick K is done with tests
                 21:10 KTEVANA errors, ordinal number mismatch, KFREND
                       KTEVANA errors again, power cycle laser crate repeater
                 21:20 Pipeline errors, reset pipeline
15397 22:15 KTEVANA errors again, Rick suspects laser crate,
                       cycle power yet again
                 23:00 Rick K has installed ET-board calib trigger for
                       KAMI tests, see log book XXIII, pg 182 for details,
                       including how to kill it if it creates trouble.

                 23:10 New estimate for beam is 3AM

                 23:30 RickT and Marj make access to look at laser crate
                        repeater for problems...but we didn't see anything
                        obvious
                 00:00 Hand over to owl shift
========================================================================
Items to Follow up on Next Shift

       Mike A. has installed new BA voltages for
       muon run, need to return to nominal after muon run.

        



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:39:03 CST